Find the value of a and b, if

Limx→∞ [x(1 + acosx) - b sin x]/x3 = 1


1
Expert's answer
2021-06-15T13:38:00-0400

"\lim_{x \to \infty}\frac{x(1 + acos(x)) - b sin (x)}{x^3}" ("\frac{\infty}{\infty}") form

since above form is an indeterminant form. therefore using L'Hopital rule


"\implies" "\lim_{x \to \infty}\frac{\frac{d}{dx}[x(1 + acos(x) - b sin(x)]}{\frac{d}{dx}x^3}"

"\implies" "\lim_{x \to \infty}\frac{\frac{d}{dx}(x + axcos(x)) - \frac{d}{dx}b sin(x)]}{\frac{d}{dx}x^3}"

"\implies" "\lim_{x \to \infty}\frac{(1 + acos(x)-axsin(x)) -b cos(x)}{3x^2}" ("\frac{\infty}{\infty}") form


again applying L'Hopital rule ,


"\implies" "\lim_{x \to \infty}\frac{\frac{d}{dx}[(1 + acos(x)-axsin(x)) -b cos(x)]}{\frac{d}{dx}3x^2}"

"\implies" "\lim_{x \to \infty}\frac{[(0 - asin(x)-axcos(x)-asin(x)) + b sin(x)]}{6x}"

"\implies" "\lim_{x \to \infty}\frac{(asin(x)-axcos(x)-asin(x) + b sin(x)}{6x}" ("\frac{\infty}{\infty}") form


again applying L'Hopital rule,


"\implies" "\lim_{x \to \infty}\frac{\frac{d}{dx}(asin(x)-axcos(x)-asin(x) + b sin(x)}{\frac{d}{dx}6x}"

"\implies" "\lim_{x \to \infty}\frac{(acos(x)+axsin(x)-acos(x)-acos(x) + b cos(x)}{6}" "\to" (1)

we get , "\lim_{x \to \infty}\frac{x(1 + acos(x)) - b sin (x)}{x^3}" "\neq" 1 for any a and b because each term of numerator contain either cos(x) or sin(x) which always oscillate between 1 and -1.

if we take a = 0 in (1) then it becoms "\frac{b cos(x) }{6}" which can never tends to 1 as x "\to" "\infty"


Hence, the given question is not Right .